OpenStudy (anonymous):

You stand on a cliff 30. m high. You throw a rock off the cliff straight up at 10. m/s. Calculate the time of flight and the final velocity of the rock as it strikes the ground at the bottom of the cliff.

OpenStudy (gebooors):

First you have to calculate rising time (t) , Then you calculate, how high stone rises (y). Then stone is in free wall. if V0 = 10 m/s, then in highest point v = 0 you can start with v = V0 - gt, solve time then y = V0t - 0,5gt^2, y is height from the top of cliff.